OpenStudy (anonymous):

How would I solve 8=√2^4x-8?

OpenStudy (wolf1728):

Is everything on the right side a square root? Does the equation read 8= Square root (2^4x-8) ?

OpenStudy (wolf1728):

Assuming the equation is 8= Square root (2^4x-8) then we square both sides: 64 = 2^4x -8 72 = 2^4x taking logs of both sides log (72) = 4x * log (2) 1.8573324964 = 4x * 0.3010299957 6.1699250014 = 4x x = 1.5424812504

OpenStudy (anonymous):

Thanks.